The area of a rectangle is 54 square feet. If the width is 2/3 of the length. What is the perimeter of the rectangle ?

Answer by richwmiller(17219)

area is found by multiplying length by width
LW=A
W=2/3L
54 ft^2=LW=L*(2/3*L)=(2/3)*(L^2)
54*(3/2)=L^2
divide 54 by 2
27*3=L^2
81=L^2
9=L
W=(2/3)*9=6
Perimeter=2L+2W
substitute 9 for L and 6 for W